OpenStudy (anonymous):

How many moles of bromine atoms are in a sample of 2.03 times 10^24 atoms

OpenStudy (anonymous):

The Avogadro number is 6.022 x 10^23. So no of moles = No of atoms / 6.022 x 10^23 = 2.03 x 10^24 / 6.022 x 10^23 = 3.37 Moles.